1 The Slovak language

The file slovak.dtx^{1} defines all the language-specific macros for the Slovak language.

For this language the macro \q is defined. It is used with the letters (t, d, l, and L) and adds a ' to them to simulate a ‘hook’ that should be there. The result looks like t’.

The macro \LdfInit takes care of preventing that this file is loaded more than once, checking the category code of the @ sign, etc.

1 ⟨∗code⟩

 2 \LdfInit{slovak}\captionsslovak

When this file is read as an option, i.e. by the \usepackage command, slovak will be an ‘unknown’ language in which case we have to make it known. So we check for the existence of \l@slovak to see whether we have to do something here.

 3 \ifx\l@slovak\@undefined
 4  \@nopatterns{Slovak}
 5  \adddialect\l@slovak0\fi

The next step consists of defining commands to switch to (and from) the Slovak language.

\captionsslovak The macro \captionsslovak defines all strings used in the four standard documentclasses provided with L^{A}TEX.

 6 \addto\captionsslovak{%
 7 \def\prefacename{\'Uvod}%
 8 \def\refname{Referencia}%
 9 \def\abstractname{Abstrakt}%
10 \def\bibname{Literat\'ura}%
11 \def\chaptername{Kapitola}%
12 \def\appendixname{Dodatok}%
13 \def\contentsname{Obsah}%
14 \def\listfigurename{Zoznam obr\'azkov}%
15 \def\listtablename{Zoznam tabuliek}%
16 \def\indexname{Index}%
17 \def\figurename{Obr\'azok}%
18 \def\tablename{Tabu\q lka}%%% special letter l with hook
19 \def\partname{\v{C}as\q t}%%% special letter t with hook
20 \def\enclname{Pr\'{\i}loha}%
21 \def\ccname{CC}%
22 \def\headtoname{Komu}%
23 \def\pagename{Strana}%
24 \def\seename{vi\q d}%%%  Special letter d with hook
25 \def\alsoname{vi\q d tie\v z}%%%  Special letter d with hook
26 \def\proofname{Proof}%  <-- needs translation

27 }

\dateslovak The macro \dateslovak redefines the command \today to produce Slovak dates.

28 \def\dateslovak{%
29 \def\today{\number\day.~\ifcase\month\or
30 janu\'ara\or febru\'ara\or marca\or apr\'{\i}la\or m\'aja\or j\'una\or

1The file described in this section has version number v1.2i and was last revised on 1996/12/23.

It was written by Jana Chlebikova (chlebik@euromath.dk).

31 j\'ula\or augusat\or septembra\or okt\'obra\or
32 novembra\or decembra\fi
33  \space \number\year}}

\extrasslovak \noextrasslovak The macro \extrasslovak will perform all the extra definitions needed for the Slovak language. The macro \noextrasslovak is used to cancel the actions of \extrasslovak. This currently means saving the meaning of one one-letter control sequence before defining it.

34 \addto\extrasslovak{\babel@save\q\let\q\v}

The slovak hyphenation patterns should be used with \lefthyphenmin set to 2

and \righthyphenmin set to 2.
35 \def\slovakhyphenmins{\tw@\tw@}

The macro \ldf@finish takes care of looking for a configuration file, setting the main language to be switched on at \begin{document} and resetting the category code of @ to its original value.

36 \ldf@finish{slovak}

37 ⟨/code⟩